OpenStudy (anonymous):

The organic compound di-n-butyl phthalate, C16H22O4(l), is sometimes used as a low-density (1.046 g·mL–1) manometer fluid. Compute the pressure (in torr) of a gas that supports a 535-mm column of di-n-butyl phthalate. The density of mercury is 13.53 g·mL–1. The organic compound di-n-butyl phthalate, C16H22O4(l), is sometimes used as a low-density (1.046 g·mL–1) manometer fluid. Compute the pressure (in torr) of a gas that supports a 535-mm column of di-n-butyl phthalate. The density of mercury is 13.53 g·mL–1. @Chemistry

OpenStudy (anonymous):

yupp

OpenStudy (aaronq):

is this at STP?

OpenStudy (aaronq):

ps. why did you write the question twice? lol anyway lets assume STP: use PV=nRT we can convert it into P=(mRT/MV) then P=dRT/M so at STP: T= 25 celsius which equals 298.73 Kelvin d= 1.046 g/mL = 0.001046 g/L (we need to use litres in the ideal gas equation) R =0.08206 Latm/Kmol M= molar mass of C16H22O4, which is (12)16+(1)22+(16)4 = 278 g/mol so plug this into: P=dRT/M once you found the pressure, P, you need to convert it into torr (because it will be in atm (atmospheres).. since 1 atm = 760 torr = 760 mm Hg multiply P by 760 and you're done. You;ll have the pressure in torr